Thando invested R10 000 in a special savings account on 15 May at an interest rate of 15% per year, compounded every three months for seven months. Interest is calculated on 1 January, 1 April, 1 July and 1 October of every year.

If fractional compounding is used for the full term of seven months, the total amount of interest received is

[1] R892,79.

[2] R898,43.

[3] R901,73.

[4] R894,04.

[5] none of the above


1
Expert's answer
2021-05-24T15:30:45-0400

Compounded interest

Formula= "Principal\times Rate\times Time"

Within first 3 months:

"10,000\times\frac{15}{100}\times\frac{2.5}{12}=312.50"


Following 3 months:

"10,312.50\times\frac{15}{100}\times\frac{3}{12}=386.71875"

Last 3 months:

"10,699.21875\times\frac{15}{100}\times\frac{1.5}{12}=200.6103516"


Total interest=312.5+386.71875+200.6103516=899.8291


None of the above